OpenStudy (anonymous):

Medal &Fan Describe the vertical asymptote and hole for the graph of x^2+x-6/x^2-9. A.Asymptote:x=2;hole x=-3 B.Asymptote:x=3;hole x=2 C.Asymptote:x=-3;hole x=3 D.Asymptote:x=3;hole x=-3

OpenStudy (ikram002p):

x^2+x-6/x^2-9 =0 (x-3)(x+2) /(x-3)(x+3)=0 (x+2)/(x+3) =0 .............( x +3 )dnt=0 ok ? x+2=0 , x=-2 got it untill nw ?

OpenStudy (anonymous):

I think so ,yeahh

OpenStudy (anonymous):

What do i do next?

OpenStudy (ikram002p):

( x +3 )dnt=0 ok when x+3=0 , x=-3 (it called vertical asymptote )

OpenStudy (anonymous):

what about the hole part?

OpenStudy (ikram002p):

nw from this one (x-3)(x+2) /(x-3)(x+3)=0 , u can see that we can cancelate (x-3) right ?

OpenStudy (ikram002p):

so the limite we cancelate x-3=0 x=3 got it ??

OpenStudy (anonymous):

Yess. Thank you so much!, i really appreciate it

OpenStudy (ikram002p):

good luck ^^

OpenStudy (anonymous):

Thanks (:

OpenStudy (ikram002p):

np :)